A girl is twice as old as her sister. Four year hence, the product of their ages will be 160. find their present ages
Answer by ptaylor(2198)


Let x = sisters present age; 4 yrs hence her age will be x+4
Then 2x= other girls present age; 4 years hence her age will be 2x+4
Sooooo
(x+4)(2x+4)=160
Using FOIL we have 2x^2+4x+8x+16=160
2x^2+12x+16=160. divide each term by 2
x^2+6x+8=80 or
x^2+6x-72=0 quadratic in standard form and it can be factored
(x+12)(x-6)=0
x=-12-----no good can't have negative age
x=6-----answer
sisters age is 6
girls age is 12
CK
(6+4)(12+4)=160
160=160
hope this helps----ptaylor